OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = 1,024 and a4 = -16? (3 points) 1 -0.25 -1 0.25

OpenStudy (anonymous):

@Hero @ganeshie8 @phi Please Help

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

\(a_1 = a_1\) \(a_2 = a_1 \cdot r\) \(a_3 = a_2 \cdot r = a_1 \cdot r^2\) \(a_4 = a_3 \cdot r = a_1 \cdot r^3\) \(...\) \(a_n = a_1r^{n - 1} \) Use the last formula with the given information to find r. Then use the last formula again with the r you found to find \(a_6\).